Examine the following series for convergence


∞Σn=1 [(n-1/(2n+3)]^n

1
Expert's answer
2022-01-31T15:02:34-0500

Solution:

Let "a_n=(\dfrac{n-1}{2n+3})^n"

Now, consider "L=\lim_{n\rightarrow \infty} {a_n}^{1/n}"

"\Rightarrow L=\lim_{n\rightarrow \infty} [{(\dfrac{n-1}{2n+3})^n}]^{1/n}
\\ \Rightarrow L=\lim_{n\rightarrow \infty} (\dfrac{n-1}{2n+3})
\\ \Rightarrow L=\lim_{n\rightarrow \infty} (\dfrac{1-\frac1n}{2+\frac3n})
\\ \Rightarrow L= \dfrac12"

Since, "L<1", so by root test, "\Sigma a_n" converges.